OpenStudy (anonymous):

what is 2w*2l=72?? trying to solve for w and l

OpenStudy (cwrw238):

you have two unknowns and only one equation so there are infinitely many solutions. you need two equations to get definite roots

OpenStudy (anonymous):

It looks like the perimeter of a rectangle is 72 and you are solving for the length and width, 4wl=72 => wl = 18. Without knowing either the length or the width you can't solve for either. (from before) You could then solve for each in terms of the other: w = 18/l l = 18/w but that's as far as you could go without knowing either one.